OpenStudy (anonymous):

The highway mileage m in miles per gallon for a compact car is approximated by m(s) = -0.025s^2 + 2.45s - 30, where s is the speed in miles per hour. What is the maximum mileage for this compact car to the nearest tenth of a mile per gallon? __________ mi/gal.

OpenStudy (anonymous):

max is at \(x=-\frac{b}{2a}=-\frac{2.45}{2\times (-.025)}\) whatever that is

OpenStudy (anonymous):

the maximum value of a quadratic equation and the answer to your problem is the vertex. to be more precise the y value of the vertex. first find the x value of the vertex using -b/2a. next find the y value by plugin it back into the equation(for you x's)